A spring of mass, m, oscillates with a frequency, f. Find an expression for the angular velocity, w. Also find the solution in form of x(t)
1
Expert's answer
2020-12-23T10:25:26-0500

By definition, the angular velocity is:


"\omega = 2\pi f"

where "f" is the frequency of the spring.

The solution is (simple harmonic motion):


"x(t) = A\cos (\omega t + \phi)"

where "A" is the magnitude of the oscillations, "\omega" is angular velocity, and "\phi" is the initial phase shift.


Answer. 1) "\omega = 2\pi f", 2) "x(t) = A\cos (\omega t + \phi)".